OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2) 17 −17 27 −27 Solve: -1/2 x + 3 = −x + 7 1/2 8 −8 2

OpenStudy (anonymous):

For the first one, 5(x) is 5x 5(8) is 40. That gives you 5x+40 on the left side. 7(x) is 7x and 7(-2) is -14.. That gives you 7x+ -14 on the right side. Add 7x to both sides. That gives you 12x+40 = -14. Now subtract 40 from both sides. 12x=-54, x is -4.5.

OpenStudy (anonymous):

The second one: Subtract 3 from both sides. You're left with -1/2x=-x+4. Now add x to both sides. 1/2x=4. You can multiply both sides by 2/1 to make x a whole number, which makes the whole equation 1x=8, which means x is 8.

OpenStudy (anonymous):

Ok well thats not one of my choices so is there a differnt way to solve it?

OpenStudy (anonymous):

@KCB I apologize, I was trying to work it in my head. I will resolve it right now.

OpenStudy (anonymous):

For the first one, you'll have to multiply out the parentheses. That's the only way to get x on one side by itself ;)

OpenStudy (anonymous):

I apologize. Number 1 is positive 27. 5x+40=7x-14 Subtract 5x from both sides. 40=2x-14 Add 14 to both sides. 54=2x x=27

OpenStudy (anonymous):

Your fine im jus tryna learn how to solve it

OpenStudy (anonymous):

@KCB (:

OpenStudy (anonymous):

So here's how you go about it step by step: 5(x + 8) = 7(x − 2) (=) 5x+40 = 7x-14 Then you can add or subtract to both sides of the equation to get X on one side by itself: (=) 40 = 7x - 5x - 14 (=) 40+14 = 7x-5x (=) 54 = 2x (=) 27 = x

OpenStudy (anonymous):

thank you . how would i solve the 1/2 one?

OpenStudy (anonymous):

The second one is similar. Start by subtracting 3 from both sides of the equation, then add 1x, and you've got all the products of x on one side. Then you can divide/multiply to solve for x.

OpenStudy (anonymous):

ok lemme try it

OpenStudy (anonymous):

Sorry msloo - you got it right, of course, just trying to show how the equations change with each step :)

OpenStudy (anonymous):

what do i add 1x to?

OpenStudy (anonymous):

Both sides of the equation. Anything you do you have to do to both sides of the equation, for the equation to still be true. So, after subtracting 3 from both sides: -1/2x + 3 = −x + 7 (=) -1/2x = -x + 7 -3 If you add 1x to both sides, you get this: (=) -1/2x + x = -x + x + 7 -3 -x + x equals zero for any x, so this is the same as (=) -1/2x + x = 7 -3 So by adding 1x to both sides, you've effectively eliminated all instances of the variabe on the right side of the equation.

OpenStudy (anonymous):

i got 2?

OpenStudy (anonymous):

@opiesche haha, it's fine

OpenStudy (anonymous):

No - but I see where it went wrong, just a small mistake at the last step. Continuing, if we have -1/2x + x = 7 -3 that's equivalent to (=) 1/2x = 4 So to get x on the left side, we have to multiply by 2 (not divide): (=) x = 4*2 So the result is 8.

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

Yea now i do its jus the 1/2 part that confused me Thank you :)